# High-dimensional quantum XYZ product codes for biased noise

### v1

@@ Abstract
Quantum XYZ product can construct a class of non-CSS codes by using three classical codes. However, before this work, their error-correcting performance is not studied in depth and whether this code construction method can be generalized to higher dimension is an open question. In this paper, we first study the error-correcting performance of the 3D Chamon code, which can be seen as a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes. Second, we show that XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which can be seen as a variant of 4D homological product and constructs a class of non-CSS codes by using 4 classical codes or 2 CSS codes. Compared with 4D homological product, we show that 4D XYZ product can construct non-CSS codes with higher dimension or code distance. Third, we consider two special instances of 4D XYZ product, which we name 4D Chamon code and 4D XYZ concatenated code. Exploiting fully decoupled binary belief propagation combined with ordered statistics decoding, our simulation results show that, using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@ I Introduction
The motivation of our work is to answer the problems above. In this paper, we first study the error-correcting performance of three-dimensional (3D) Chamon codes[14, 15], which is a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes with length $n_{1}$, $n_{2}$, $n_{3}$. Exploiting fully decoupled binary belief propagation combined with ordered statistics decoding (FDBP-OSD[16]), when $n_{1}=n_{2}=n_{3}$, the code-capacity threshold of the corresponding Chamon codes is around $14\%$ under depolarizing noise model. For pure Pauli $X$, $Y$ and $Z$ noise model, the corresponding code-capacity threshold of Chamon codes is around $13\%$, $14\%$, $14\%$, respectively. Second, we show that that XYZ product can be generalized to four dimension and propose 4D XYZ product code construction, which can be seen as a variant of standard homological product of two CSS codes (namely, 4D homological product) and constructs a non-CSS code by making use of 4 classical codes or 2 CSS codes. Third, we study the dimension and code distance of 4D XYZ product codes, and show that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product. Finally, we explore the error-correcting performance of two special instances of 4D XYZ product. The first one is constructed from two quantum concatenated codes, which are obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$) (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd). The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$). We name these two codes 4D XYZ concatenated code and 4D Chamon code, respectively. Exploiting FDBP-OSD, we study their error-correcting performances under the independent single-qubit Pauli noise model. Our simulation results show that under pure Pauli $Z$ noise the code-capacity thresholds of 4D XYZ concatenated codes and 4D Chamon codes are much higher than those of the 4D toric codes[17] and 4D homological concatenated codes (namely, the codes constructed by 4D homological product of two quantum concatenated codes, which are the same two codes used to construct 4D XYZ concatenated code) respectively, while their code-capacity thresholds are close under depolarizing noise. Besides, under $Z$-biased noise model with different bias rate, from the perspective of code-capacity threshold, 4D XYZ concatenated codes and 4D Chamon codes also show good error-correcting performance. Our simulation results means that, using the same two codes, 4D XYZ product can constructed non-CSS codes with better error-correcting performance for biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@ IV Numerical simulations
In this section, we exploit FDBP-OSD proposed in Ref. [16] to study the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes under the independent single-qubit Pauli noise model, in which each qubit independently suffers a Pauli error $I$, $X$, $Y$, or $Z$ with error probability $1-p$, $p_{x}$, $p_{y}$, $p_{z}$, respectively, where $p=p_{x}+p_{y}+p_{z}$ is physical qubit error rate. Under this noise model, the bias rate of Pauli $Z$ noise is defined as
@@
We should emphasize that, in this paper we do not explore an efficient and highly accurate decoder for 3D and 4D XYZ product codes, and FDBP we used here is an improved binary BP, which can be applied to all quantum stabilizer codes and have better decoding performance than tradition binary BP[21, 22], especially for some non-CSS codes[16]. However, as shown in Ref. [23], even though tradition binary BP-OSD[24] has been shown to have pretty good decoding performance for a wide range of QLDPC codes, but its performance can degrade for the codes with large code length and high degeneracy. FDBP also has similar problem and its decoding performance is also greatly degraded by short cycles, especially 4-cycles in the Tanner graph. In Appendix B, we compare the number of 4-cycles in the Tanner graphs of 3D Chamon code, 3D toric codes, 4D Chamon codes and 4D toric codes. Our computed result shows that the number of 4-cycles in the Tanner graph of 3D XYZ and 4D XYZ product codes is much higher than that of 3D and 4D homological product codes. Thus, the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes obtained by FDBP-OSD in this paper is not the optimal. Nonetheless, from the perspective of code-capacity threshold, they still show good error-correcting performance for $Z$-biased noise model.
@@
In Fig. 8, we first consider the case of $n_{1}=n_{2}=n_{3}=n_{4}$. Fig. 8(a) and (b) show the single logical qubit error rate against the physical qubit error rate $p$ of 4D Chamon codes and 4D toric codes under depolarizing noise model, respectively. In this case, the 4D Chamon code encodes $8n_{1}^{2}$ logical qubits, while the 4D toric codes always encodes 6 logical qubits. In the legend, we use notation $L=n_{1},n_{2},n_{3},n_{4}$ to represent the corresponding the code, which is constructed by four repetition codes with code length of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. One can see that the code-capacity threshold of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD is around $12.8\%$ and $15.9\%$, respectively, and the single logical qubit error rate of 4D Chamon codes is much higher than that of 4D toric codes. We attribute this to the fact that the number of 4-cycles of 4D Chamon codes is much more than that of 4D toric codes as shown in Appendix B, which degrades the decoding performance of FDBP-OSD.
@@
The dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, when $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the corresponding code always encode 8 logical qubits, which is close to 4D toric codes. Thus, it is more reasonable to compare the error-correcting performance between 4D Chamon codes and 4D toric codes in this case. First, we consider the pure Pauli $Z$ noise model. As shown in Fig. 9 (a) and (b), the code capacity of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD under Pauli $Z$ noise is around $18\%$ and $9.5\%$, respectively. This result means, form the perspective of code-capacity threshold, 4D Chamon codes have better error-correcting performance than that of 4D toric codes under pure Pauli $Z$ noise.
@@
Figure 11: The error-correcting performance of (a) 4D XYZ concatenated code and (b) 4D homological product codes under pure $Z$ noise model.
@@ Appendix B Comparison of number of 4-cycles
Table 2 gives the number of 4-cycles in the Tanner graphs of 3D Chamon codes, 4D Chamon codes, 3D toric codes and 4D toric codes.

### v2

@@ Abstract
Quantum XYZ product can construct a class of non-CSS codes by using three classical codes. However, before this work, their error-correcting performance is not studied in depth and whether this code construction method can be generalized to higher dimension is an open question. In this paper, we first study the error-correcting performance of the 3D Chamon code, which can be seen as a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes. Second, we show that XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which can be seen as a variant of 4D homological product and constructs a class of non-CSS codes by using 4 classical codes or 2 CSS codes. Compared with 4D homological product, we show that 4D XYZ product can construct non-CSS codes with higher dimension or code distance. Third, we consider two special instances of 4D XYZ product, which we name 4D Chamon code and 4D XYZ concatenated code. Exploiting fully decoupled binary belief propagation combined with $order\mbox{-}0$ ordered statistics decoding, our simulation results show that, using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@ I Introduction
The motivation of our work is to answer the problems above. In this paper, we first study the error-correcting performance of three-dimensional (3D) Chamon codes[14, 15], which is a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes with length $n_{1}$, $n_{2}$, $n_{3}$. Exploiting fully decoupled binary belief propagation combined with $order\mbox{-}0$ ordered statistics decoding (FDBP-OSD-0[16]), when $n_{1}=n_{2}=n_{3}$, the code-capacity threshold of the corresponding Chamon codes is around $14\%$ under depolarizing noise model. For pure Pauli $X$, $Y$ and $Z$ noise model, the corresponding code-capacity threshold of Chamon codes is around $13\%$, $14\%$, $14\%$, respectively. Second, we show that that XYZ product can be generalized to four dimension and propose 4D XYZ product code construction, which can be seen as a variant of standard homological product of two CSS codes (namely, 4D homological product) and constructs a non-CSS code by making use of 4 classical codes or 2 CSS codes. Third, we study the dimension and code distance of 4D XYZ product codes, and show that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product. Finally, we explore the error-correcting performance of two special instances of 4D XYZ product. The first one is constructed from two quantum concatenated codes, which are obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$) (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd). The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$). We name these two codes 4D XYZ concatenated code and 4D Chamon code, respectively. Exploiting FDBP-OSD-0, we study their error-correcting performances under the independent single-qubit Pauli noise model. Our simulation results show that under pure Pauli $Z$ noise the code-capacity thresholds of 4D XYZ concatenated codes and 4D Chamon codes are much higher than those of the 4D toric codes[17] and 4D homological concatenated codes (namely, the codes constructed by 4D homological product of two quantum concatenated codes, which are the same two codes used to construct 4D XYZ concatenated code) respectively, while their code-capacity thresholds are close under depolarizing noise. Besides, under $Z$-biased noise model with different bias rate, from the perspective of code-capacity threshold, 4D XYZ concatenated codes and 4D Chamon codes also show good error-correcting performance. Our simulation results means that, using the same two codes, 4D XYZ product can constructed non-CSS codes with better error-correcting performance for biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@ IV Numerical simulations
In this section, we exploit FDBP-OSD-0 proposed in Ref. [16] to study the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes under the independent single-qubit Pauli noise model, in which each qubit independently suffers a Pauli error $I$, $X$, $Y$, or $Z$ with error probability $1-p$, $p_{x}$, $p_{y}$, $p_{z}$, respectively, where $p=p_{x}+p_{y}+p_{z}$ is physical qubit error rate. Under this noise model, the bias rate of Pauli $Z$ noise is defined as
@@
We should emphasize that, in this paper we do not explore an efficient and highly accurate decoder for 3D and 4D XYZ product codes, and FDBP we used here is an improved binary BP, which can be applied to all quantum stabilizer codes and have better decoding performance than tradition binary BP[21, 22], especially for some non-CSS codes[16]. However, as shown in Ref. [23], even though tradition binary BP-OSD[24] has been shown to have pretty good decoding performance for a wide range of QLDPC codes, but its performance can degrade for the codes with large code length and high degeneracy. FDBP also has similar problem and its decoding performance is also greatly degraded by short cycles, especially 4-cycles in the Tanner graph. In Appendix B, we compare the number of 4-cycles in the Tanner graphs of 3D Chamon code, 3D toric codes, 4D Chamon codes and 4D toric codes. Our computed result shows that the numbers of 4-cycles in the Tanner graph of 3D and 4D XYZ product codes are much more than those of 3D and 4D homological product codes. Thus, the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes obtained by FDBP-OSD-0 in this paper is not the optimal. Nonetheless, from the perspective of code-capacity threshold, they still show good error-correcting performance for $Z$-biased noise model.
@@
In Fig. 8, we first consider the case of $n_{1}=n_{2}=n_{3}=n_{4}$. Fig. 8(a) and (b) show the single logical qubit error rate against the physical qubit error rate $p$ of 4D Chamon codes and 4D toric codes under depolarizing noise model, respectively. In this case, the 4D Chamon code encodes $8n_{1}^{2}$ logical qubits, while the 4D toric codes always encodes 6 logical qubits. In the legend, we use notation $L=n_{1},n_{2},n_{3},n_{4}$ to represent the corresponding code, which is constructed by four repetition codes with code length of $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$. One can see that the code-capacity threshold of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 is around $12.8\%$ and $15.9\%$, respectively, and the single logical qubit error rate of 4D Chamon codes is much higher than that of 4D toric codes. We attribute this to the fact that the number of 4-cycles of 4D Chamon codes is much more than that of 4D toric codes as shown in Appendix B, which degrades the decoding performance of FDBP-OSD-0.
@@
The dimension of 4D Chamon code is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$, when $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are coprime, the corresponding code always encode 8 logical qubits, which is close to 4D toric codes. Thus, it is more reasonable to compare the error-correcting performance between 4D Chamon codes and 4D toric codes in this case. First, we consider the pure Pauli $Z$ noise model. As shown in Fig. 9 (a) and (b), the code capacity of 4D Chamon codes and 4D toric codes obtained by FDBP-OSD-0 under Pauli $Z$ noise is around $18\%$ and $9.5\%$, respectively. This result means, form the perspective of code-capacity threshold, 4D Chamon codes have better error-correcting performance than that of 4D toric codes under pure Pauli $Z$ noise.
@@
Figure 11: The error-correcting performance of (a) 4D XYZ concatenated codes and (b) 4D homological concatenated codes under pure $Z$ noise model.
@@ Appendix B Comparison of number of 4-cycles
Table 2 gives the number of 4-cycles in the Tanner graphs of 3D Chamon codes, 4D Chamon codes, 3D toric codes and 4D toric codes. One can see that the numbers of 4-cycles in the Tanner graph of 3D and 4D Chamon codes are much higher than those of 3D and 4D toric codes.